The mean age of a class of 25 students is 23.8 years. How old would a 26th student have to be for the mean age of the class to be 24.5 years?

Answer by mananth(16946)

The mean age of a class of 25 students is 23.8 years. How old would a 26th student have to be for the mean age of the class to be 24.5 years?
sum of the ages of 25 students = 25*23.8= 595
..
one student added so total 26 students
average required = 24.5
total of ages = 26*24.5 =637
..
637-595 = 42 years should be his age
...
m.ananth@hotmail.ca